A compound contains 32%carbon,4%hydrogen, and rest oxygen.its vapour density is 75.calculate the empirical and molecular formula.

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
For computation the emperical and sub-atomic equation initial step is discover the no gram molecules of carbon and hydrogen. 
No of gram particles of carbon= 32/12= 2.66 
No of gram particles of hydrogen=4/1= 4 
No of gram particles of rest oxygen=63.16/16= 3.94 
Nuclear proportion of C:H:O 2.66:4:3.94 
All the nuclear proportion is partitioned into littlest no gram molecules 
C:H:O 2.66/2.66:4:2.66:3.94/2.66= 1:1.50:1.40 
To change over the part into entire no appropriate digit 2 is duplicated with nuclear proportion 
C:H:O 2(1:1.50:1.40)= 2:3:3 
The observational equation of the compound is C2H3O3 
The sub-atomic equation is C2H3O3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)